OpenStudy (anonymous):

Choose the missing number to make equivalent fractions. 7/a = 24/40

OpenStudy (anonymous):

@FuyukoHarayo Want another medal?

OpenStudy (anonymous):

A. 10 B. 19 C. 28 D. 160

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@FuyukoHarayo It is ok if you can't answer.

OpenStudy (anonymous):

This is kinda hard. >-<;

OpenStudy (anonymous):

It is ok.

OpenStudy (anonymous):

@lizzyrockefeller @TheSmartOne Want an extra medal?

OpenStudy (danjs):

do a cross multiply first

OpenStudy (anonymous):

il draw it out

OpenStudy (danjs):

\[\frac{ 7 }{ a } = \frac{ 24 }{ 40 }\] 7*40 = a*24

OpenStudy (danjs):

that enough to finish it?

OpenStudy (anonymous):

I'm not possative.

OpenStudy (danjs):

cross multiply gives 7*40 = 24a 280 = 24a divide both sides by 24 a = 280/24 = 35/3 none of the answers are that

OpenStudy (danjs):

\[\frac{ 7 }{ a } = \frac{ 24 }{ 40 }\] That the correct problem?

OpenStudy (anonymous):